Apologies for asking a basic question;
A triangle, one angle 85, other X, and third X-7
Is it possible to provide an equation to work this out?
Many thanks,

Answer by ankor@dixie-net.com(22740)

A triangle, one angle 85, other X, and third X-7
:
A simple equation
x + (x-7) + 85 = 180
2x + 78 = 180
2x = 180 - 78
2x = 102
x = 102/2
x = 51
:
Therefore the 2nd and 3rd angles are 51 and 44 degrees
:
Check: 44 + 51 + 85 = 180